### v1

Abstract
This work presents the results of a methodological transfer from remote sensing to healthcare, adapting AMBER (Dosi et al., 2025) — a transformer-based model originally designed for multiband images such as hyperspectral data — to the task of 3D medical datacube segmentation. In this study, we use the AMBER architecture with Adaptive Fourier Neural Operators (AFNO) in place of the multi-head self-attention mechanism. While existing models rely on various forms of attention to capture global context, AMBER-AFNO achieves this through frequency-domain mixing, enabling a drastic reduction in model complexity. This design reduces the number of trainable parameters by over 80% compared to UNETR++, while maintaining a FLOPs count comparable to other state-of-the-art architectures. Model performance is evaluated on two benchmark 3D medical datasets—ACDC and Synapse—using standard metrics such as Dice Similarity Coefficient (DSC) and Hausdorff Distance (HD), demonstrating that AMBER-AFNO achieves competitive or superior accuracy with significant gains in training efficiency, inference speed, and memory usage.
keywords:
[…]
1 Introduction
According to the World Health Statistics 2024 report (World Health Organization (2024)), heart disease was the leading cause of death globally in 2021, accounting for approximately 9.1 million deaths, with kidney disease, lung cancer, lower respiratory infections, and stroke also ranking among the top ten global causes of mortality.
In all cases, an early diagnosis is crucial to reduce mortality rates as well as to optimize the choice amongst different treatment possibilities such as different therapies or even surgery. This early diagnosis heavily depends upon on the proper exploitation of advanced technologies such as, for instance, endoscopy, MRI, and CT Scans which allow the detection of tissue abnormalities (Tai et al., 2019). MRI and CT Scans generate reports capturing slices at various depths, resulting in three-dimensional images (hereafter data cubes) that capture more information than the 2D images (b/W or RGB) produced by traditional X-ray machines Florkow et al. (2022).
As we shall describe in more detail in the next section, most healthcare datacubes represent a 3D volume of the anatomy. Since a single 2D slice alone often fails to capture the full anatomical context, relying solely on such data can compromise model performance (Zhou et al., 2019). It is therefore no surprise that, in recent years, considerable effort has been devoted to developing advanced methods, often based on artificial intelligence, that can fully exploit the richer information embedded in these volumetric datasets.
In medical image segmentation tasks, the U-Net architecture (Ronneberger et al., 2015) has gained widespread popularity (Sathianathen et al., 2022; Bakas et al., 2019) due to its encoder–decoder structure, which effectively captures both local and global contextual representations. The skip connections allow the model to recover spatial information lost during downsampling in the encoder, making U-Net well-suited for pixel-wise segmentation tasks. Building upon this foundation, a variety of U-Net-based models have been proposed to address its inherent limitations, such as fixed receptive fields and difficulty in capturing multi-scale features. Notable extensions include AFFU-Net (Zheng et al., 2022), OAU-Net (Song et al., 2023), ISTD-Net (Hou et al., 2022), MultiResU-Net (Lan et al., 2022), SAU-Net (Chen et al., 2024), KiU-Net (Valanarasu et al., 2022), and UCR-Net (Sun et al., 2022). These models preserve the encoder–decoder backbone but introduce innovations such as adaptive fusion, dual branches, attention mechanisms, and multi-resolution pathways to improve segmentation accuracy and robustness.
Despite these advances, convolutional neural networks (CNNs), including U-Net and its variants, are inherently limited in capturing long-range dependencies due to their reliance on kernel sizes and stride settings. To address this issue, researchers have proposed architectural enhancements such gradient-guided learning (e.g., LMISA (Jafari et al., 2022)) to improve global context modeling. However, these enhancements only partially overcome limitations, particularly in the context of volumetric data such as CT or MRI scans, where modeling spatial relationships across multiple slices becomes increasingly important.
To overcome these shortcomings, transformer architectures (Vaswani et al., 2017), originally developed for natural language processing, have been adapted to vision tasks through Vision Transformers (ViT) (Dosovitskiy et al., 2021). Unlike CNNs, transformers rely on self-attention mechanisms rather than convolution operations, allowing them to model long-range dependencies across the entire image. This has led to a new class of transformer-based models for medical image segmentation, including RT-UNet (Li et al., 2022a), SWTRU (Zhang et al., 2022), LMIS (Zhu et al., 2024b) and SDV-TUNet (Zhu et al., 2024a). These models often integrate transformer blocks into U-Net-like frameworks to benefit from both spatial precision and global context modeling. However, they typically require deep encoder stacks—often with 8 to 12 transformer layers—and rely on computationally expensive attention mechanisms, making them less efficient in terms of memory and inference time compared to lightweight CNN counterparts.
SegFormer (Xie et al., 2021) addresses key limitations in semantic segmentation by introducing a more efficient design, featuring a lightweight mix transformer block and a simple MLP-based decoder. It removes positional encoding in favor of a feed-forward network (FFN) and streamlines the attention mechanism for greater computational efficiency. Its encoder, composed of just four hierarchical mix transformer blocks, is significantly lighter than those in traditional architectures.
Building upon this foundation, AMBER (Dosi et al., 2025) extends SegFormer to multi-band image segmentation, incorporating three-dimensional convolutions, custom kernel sizes, and a Funnelizer layer to better handle complex spatial-spectral relationships.
In this work, we introduce AMBER-AFNO (where AFNO stands for Adaptive Fourier Neural Operators, see below), a variant of AMBER fine tuned for 3D medical datacube segmentation. This model represents a methodological transfer from remote sensing to healthcare, adapting the original AMBER—designed for multiband images such as hyperspectral data—to volumetric medical imaging.
In AMBER-AFNO, we replace the conventional multi-head self-attention mechanism with Adaptive Fourier Neural Operators (Guibas et al. (2021)), enabling global context modeling through frequency-domain mixing rather than attention. This results in a substantial reduction in architectural complexity, cutting the number of trainable parameters by over 80% compared to UNETR++, while preserving a number of FLOPs comparable to other state-of-the-art models.
We trained and validated our AMBER-AFNO approach by conducting comprehensive experiments on two benchmarks: ACDC (Bernard et al., 2018) and Synapse (Landman et al., 2015). Both the evaluation metrics: Dice score (DSC) and Hausdorff distance (HD) score, along with the model architecture in terms of total number of trainable parameter shows the efficiency of AMBER-AFNO compared to the existing methods in the literature.
2 Related Work
Three-dimensional medical-image segmentation has evolved rapidly since the success of the encoder–decoder paradigm initiated by U–Net2. Current methods can be grouped into three non-exclusive research lines that progressively trade architectural simplicity for richer multi-scale context awareness:
  1. 1.
    Single-branch encoder–decoder networks
  2. 2.
    Multi-branch encoder networks
  3. 3.
    Efficient Attention Methods
Single-branch models remain attractive for their efficiency in both training and inference. V-Net introduced residual 3-D convolutions and Dice loss to tackle volumetric data directly (Milletari et al., 2016). The DeepLab family incorporates Atrous convolutions and the Atrous Spatial Pyramid Pooling (ASPP) module to enlarge the receptive field without increasing the number of parameters (Chen et al., 2018). nnU-Net argues against over-engineered designs, proposing a self-configuring pipeline that automatically adapts its depth, patch size and learning schedule to each dataset and has become the de-facto baseline in many challenges (Isensee et al., 2020).
Multi-branch encoders explicitly separate semantic scales to improve both global context perception and boundary precision. DS-TransUNet employs dual-scale ViT encoders combined with a Token-Interaction Fusion module, achieving superior Dice on abdominal CT (Lin et al., 2022). DHR-Net couples a multi-scale branch with a detail-enhancement branch selected by reinforcement learning to boost small-structure recall (Bai et al., 2024). Further examples include MILU-Net, which mitigates information loss with dual up-sampling paths (He et al., 2024), and BSC-Net, a dual-resolution lightweight design that matches U-Net++ accuracy with roughly half the parameters (Zhou et al., 2024). Despite their merits, many two-branch schemes are hard to migrate across datasets and remain computationally heavy.
[…]
In this work, we introduce a fourth, original research direction, grounded in the use of Adaptive Fourier Neural Operators (AFNO)—a novel and highly efficient alternative to dense self-attention. Unlike traditional attention, AFNO leverages spectral-domain mixing to capture long-range dependencies with drastically fewer learnable parameters and lower memory requirements. By projecting features into the Fourier domain, performing learnable, adaptive filtering on a truncated set of low-frequency modes, and transforming back to the spatial grid, AFNO enables global context modeling at quasi-linear complexity.
This mechanism avoids the quadratic scaling of token–token attention and offers excellent scalability to 3D medical datacubes, making it especially attractive for segmentation tasks in clinical settings where memory and compute resources may be limited. To the best of our knowledge, this is the first time AFNO has been applied to volumetric medical imaging tasks, thus establishing a new, efficient paradigm for transformer-based 3D segmentation.
When integrated into hierarchical encoders, AFNO modules serve as a drop-in alternative to windowed or factorized attention schemes (e.g., Swin, EPA, G-SWPA), offering a different efficiency trade-off: FLOPs remain comparable to conventional transformer stages, but parameter counts shrink substantially because mixing weights operate over frequency bands rather than full token embeddings (see tab. 3 and 5). Additionally, the ability of AFNO modules to take advantage of frequency-based mixing enhances their performance on high-dimensional data typical in medical imaging, where capturing fine-grained details is crucial
In summary, the evolution from streamlined single-branch CNNs to increasingly elaborate multi-branch and attention-centric transformers has steadily enhanced volumetric segmentation performance, albeit at the cost of growing model bulk and deployment friction. Our AFNO-based variant, AMBER-AFNO, recovers the global-context benefits of self-attention while replacing token-token interactions with light-weight spectral mixing, reducing the number of parameters without reducing performance.
3 Methodology
In this section, we introduce AMBER-AFNO, an extension of the AMBER model (Dosi et al., 2025), which replaces the traditional self-attention mechanism with Adaptive Fourier Neural Operators (AFNO) to reduce model complexity and improve computational efficiency. While staying close to the original AMBER design, AMBER-AFNO adapts the architecture to address the 3D volumetric data. As illustrated in Figure 1, the architecture consists of two main modules: a hierarchical Transformer encoder with 3D patch embedding and AFNO-based feature mixing; and a lightweight MLP decoder that fuses multi-scale features and predicts the final 3D segmentation mask. Unlike AMBER, which uses a dimensionality reduction layer (“funnalizer”) to collapse 3D features into 2D outputs, AMBER-AFNO operates entirely in 3D and directly outputs a volumetric segmentation mask of size $D\times H\times W\times N_{cls}$, where $N_{cls}$ is the number of classes. Furthermore, the decoder integrates a deconvolutional layer to up-sample feature maps and recover the original spatial resolution.
Figure 1: The Proposed AMBER-AFNO framework consists of two main modules: A hierarchical Transformer encoder to extract coarse and fine features; and a lightweight MLP decoder to directly fuse these multi-level features and predict the semantic segmentation mask. FFN indicates a feed-forward network.
3.1 Hierarchical Transformer Encoder
We designed a series of Mix Transformer encoders (MiT) for semantic segmentation of 3D images, replacing standard self-attention layers with Adaptive Fourier Neural Operators (AFNO) to reduce the number of parameters without compromising accuracy. Unlike conventional attention mechanisms—which are memory-intensive and scale quadratically with input size—AFNO performs global feature mixing in the frequency domain with quasi-linear complexity. This enables the encoder to capture both local and global context while significantly lowering computational and memory costs.
Hierarchical Feature Representation. The goal of this module is, given an input image, to generate CNN-like multi-level features. These features provide high-resolution coarse features and low-resolution fine-grained features that usually boost the performance of semantic segmentation (Xie et al., 2021). More precisely, given an input 3D image with $D\times H\times W$, we perform patch merging to obtain a hierarchical feature $F_{1}$ with a resolution of $\frac{D_{qw}}{2^{i}}\times\frac{H}{2^{i}}\times\frac{W}{2^{i}}\times C_{1}$, where $i\in\{0,1,2,3\}$ and $C_{i+1}$ is larger than $C_{i}$.
Overlapped Patch Merging. We utilize merging overlapping patches to avoid the need for positional encoding. To this end, we define $K$, $S$, and $P$, where $K$ is the three-dimensional kernel size (or patch size), $S$ is the stride between two adjacent patches, and $P$ is the padding size. Unlike the original SegFormer, in our experiments, we set $K$ = 3, $S$ = 1, $P$ = 1, and $K$ = 3, $S$ = 2, $P$ = 2 to perform overlapping patch merging. The patch size is intentionally kept small to preserve image details and avoid parameter explosion. $S$ = 1 preserves the original image spatial dimensions $H$ and $W$, avoiding the reduction of spatial dimensions by 1/4.
Adaptive Fourier Neural Operators (AFNO). The self-attention mechanism, while effective in capturing global dependencies, is widely recognized as the primary computational bottleneck in transformer-based encoder architectures due to its quadratic complexity with respect to the input sequence length (Xie et al., 2021).
In AMBER (Dosi et al., 2025), a multi-head self-attention scheme is used, where each attention layer incorporates a reduction factor $R$ to mitigate this cost. This reduces the computational complexity from the standard $O(N^{2})$ to $O\left(\frac{N^{2}}{R}\right)$, offering a more tractable solution for high-resolution inputs (Xie et al., 2021).
In the proposed AMBER-AFNO architecture, the entire attention mechanism is replaced with the AFNO block. AFNO uses token mixing in the frequency domain, leveraging the Fast Fourier Transform (FFT) to achieve global interactions with quasi-linear complexity. Here, the input tokens are transformed into the frequency domain using FFT, capturing global contextual information efficiently.
In (Guibas et al., 2021) the authors introduced AFNO for the 2D image segmentation task. Using similar methodologies, we have extended the approach for the 3D image segmentation task.
The input to the AFNO block is a 5-D tensor which can be represented as $x\in\mathbb{R}^{B\times D\times H\times W\times C}$ where $B$ is the batch size, $D$, $H$, and $W$ are the spatial dimensions (depth, height, and width), and $C$ is the embedding dimension.
We then apply a real-valued 3D Fast Fourier Transform (RFFT) over the spatial dimensions:
$\hat{x}=\text{RFFT}_{3}(x)\in\mathbb{C}^{B\times D\times H\times(W/2+1)\times C}$ (1)
The channel dimension $C$ is partitioned into $K$ frequency blocks:
$\hat{x}\rightarrow\hat{x}_{\text{blk}}\in\mathbb{C}^{B\times D\times H\times(W/2+1)\times K\times\frac{C}{K}}$ (2)
Each frequency block undergoes a learnable complex-valued two-layer MLP:
$\hat{x}_{\text{blk}}^{(i)}\leftarrow W_{2}^{(i)}\cdot\phi\left(W_{1}^{(i)}\cdot\hat{x}_{\text{blk}}^{(i)}+b_{1}^{(i)}\right)+b_{2}^{(i)},\quad\forall i\in\{1,\dots,K\}$ (3)
After the operation of two layer MLP, the MLP output is reshaped back to $\hat{x}\in\mathbb{C}^{B\times D\times H\times(W/2+1)\times C}$ and then soft shrinkage is applied to attenuate small-magnitude frequency responses:
$\hat{x}^{\prime}=\text{SoftShrink}(\hat{x})$ (4)
Finally, the inverse 3D FFT (IRFFT) is applied, and the result is combined with the original input via residual addition:
$\tilde{x}=\text{IRFFT}_{3}(\hat{x}^{\prime})+x$ (5)
Thus, the final output of the AFNO-3D block is as follows:
$\boxed{\tilde{x}=\text{IRFFT}_{3}\left(\text{SoftShrink}\left(\text{MLP}_{\mathbb{C}}\left(\text{RFFT}_{3}(x)\right)\right)\right)+x}$ (6)
def AFNO3D(x):
    bias = x
    x = RFFT3(x)
    x = x.reshape(b, d, h, w//2+1, k, c/k)
    x = BlockMLP(x)
    x = x.reshape(b, d, h, w//2+1, c)
    x = SoftShrink(x)
    x = IRFFT3(x)
    return x + bias
x = Tensor[b, d, h, w, c]
W_1, W_2 = ComplexTensor[k, c/k, c/k]
b_1, b_2 = ComplexTensor[k, c/k]
def BlockMLP(x):
    x = MatMul(x, W_1) + b_1
    x = ReLU(x)
    return MatMul(x, W_2) + b_2
Figure 2: Pseudocode for AFNO-3D with adaptive weight sharing and adaptive masking
Figure 2 shows the pseudo-implementation of the AFNO-3D block in the AMBER-AFNO architecture.
Mix-FFN. Likewise, in the AMBER (Dosi et al., 2025) and SegFormer (Xie et al., 2021), we also used the Mix-FFN, which considers the effect of zero padding using a $\times 3\times 3$ 3 Conv in the feed-forward network (FFN). We used Mix-FFN instead of positional encoding because Mix-FFN combines both Depthwise Convolution and MLP layers to capture both local and global context in the seen (Xie et al., 2021).
$x_{\text{out}}=\text{MLP}\left(\text{GELU}\left(\text{Conv}_{3\times 3\times 3}\left(\text{MLP}(x_{\text{in}})\right)\right)\right)+x_{\text{in}}$ (7)
where $x_{in}$ is the feature from the self-attention module. Mix-FFN mixes a $\times 3\times 3$ 3 convolution and an MLP into each FFN.
3.2 Lightweight All-MLP Decoder
The four feature maps produced by the MiT encoder are first channel-projected to a common embedding size $d$ by position-agnostic MLP layers. Then every projected tensor is trilinearly upsampled to the finest encoder resolution and concatenated along the channel axis. A $\times 1\times 1$ 1 convolution with ReLU activation and 3-D batch normalization fuses this aggregate into a compact representation. A single transposed 3-D convolution subsequently doubles the spatial dimensions, bringing the volume back to the native voxel grid while reducing the channels to $N_{\mathrm{cls}}$. A final $\times 1\times 1$ 1 convolution sharpens the logits, yielding the segmentation mask $M\in\mathbb{R}^{N_{\mathrm{cls}}\times D\times H\times W}$. Compared with the original five-stage AMBER decoder, this variant omits the dedicated spectral-reduction block and integrates the final MLP into the transposed convolution, lowering memory cost without sacrificing accuracy.
The figure 3 shows a more simplified and straightforward workflow diagram of the AMBER-AFNO Architecture.
Figure 3: Simplified Work Flow Diagram of AMBER-AFNO Architecture
4 Experiments
To assess the effectiveness of the proposed AMBER-AFNO architecture, we performed a comprehensive benchmark against a selection of representative convolutional and transformer - based segmentation models. These include U-Net Ronneberger et al. (2015), TransUNet (Chen et al., 2021), Swin-UNet (Cao et al., 2023), UNETR (Hatamizadeh et al., 2021), MISSFormer (Huang et al., 2023), Swin-UNETR (Hatamizadeh et al., 2022), nnFormer (Zhou et al., 2023), UNETR++ (Shaker et al., 2024), LeVit-UNet (Feng et al., 2024), and PCCTrans (Xu et al., 2024). While this list is not exhaustive, it represents a diverse and competitive set of baselines from both CNN and transformer families.
All models are evaluated on two heterogeneous public benchmarks. ACDC cardiac MR challenge and the Synapse multiorgan abdominal CT segmentation task. The following subsections detail the datasets and the pre-processing protocols used, the training and inference procedures adopted throughout this study, and the quantitative criteria — Hausdorff distance (HD95) and Dice similarity coefficient (DSC) — used for performance comparison.
4.1 Dataset Overview
Automated cardiac diagnostic segmentation dataset (ACDC)
The ACDC dataset (Bernard et al., 2018) consists of 3D cardiac MRI images with multi-class annotations. A total of 200 labeled samples were used, split into 160 for training and 40 for testing. The annotated classes include the right ventricle (RV), myocardium (MYO), and left ventricle (LV). In this study, the Dice Similarity Coefficient (DSC) was used as the evaluation metric for model comparison.
Multi-organ ct segmentation dataset (Synapse)
The Synapse dataset (Landman et al., 2015) includes 30 abdominal CT scans, divided into 18 training samples and 12 evaluation samples using the UNETR++(Shaker et al., 2024) processing approach. It features eight segmentation targets: spleen, right kidney, left kidney, gallbladder, liver, stomach, aorta, and pancreas. For performance assessment, this study utilized the Dice Similarity Coefficient (DSC) and the 95th percentile Hausdorff Distance (HD95) as evaluation metrics.
Table 1: High-level description of the datasets.
Dataset Spatial Dimension Depth Dimension Modality Class Training Samples Test Samples
ACDC 320 × 320 (resampled) 90–130 slices 1 (Cine MRI) 4 160 cases 40 cases
Synapse 512 × 512 75–250 slices 1 (CT) 9 18 scans 12 scans
5 Experimental Settings
We have used the same preprocessing steps which are used such as Resampling, Intensity Normalization, Z- Score Normalisation, Cropping and Padding, in the UNETR++ (Shaker et al., 2024), UNETR (Hatamizadeh et al., 2021), DS-UNETR++ (Jiang et al., 2025), SAM (Cen et al., 2024) Models.
The program is executed on a single NVIDIA Tesla V100-SXM2 32G GPU with Thermal Design Power (TDP) of 300W. The CUDA version used is cu12.2. For all data sets, the learning rate was set to 0.01 and weight decay to 3e-5. The Deep Supervision (Li et al., 2022b) technique was used to formulate the hybrid loss function combining the Dice and Cross Entropy Loss.
6 Loss Function
For the semantic segmentation task, one possible problem that arises is class imbalance. We have considered 4 datasets with Binary, Ternary, and Quaternary segmentation problems. All the datasets have class imbalance issues. The background class dominates over other classes as the ROI for medical images is very low compared to the original shape of the image. One loss function alone, such as focal loss, dice loss or cross entropy loss can not handle this situation. To address this issue we have used a custom weighted loss function using the concept of Deep Supervision, which is a combination of cross-entropy loss and dice loss. Instead of relying on the final output of the model, the intermediate feature maps or predictions at different resolutions of the decoder block are also considered while calculating the loss and backpropagation.
$L(G,P)=1-$ $\frac{2}{J}\sum_{j=1}^{J}\frac{\sum_{i=1}^{I}G_{i,j}P_{i,j}}{\sum_{i=1}^{I}G_{i,j}^{2}+\sum_{i=1}^{I}P_{i,j}^{2}}$ (8)
$-\frac{1}{I}\sum_{i=1}^{I}\sum_{j=1}^{J}G_{i,j}\log P_{i,j}$
Where $\mathbf{G}$ refers to the set of ground truth labels, and $\mathbf{P}$ refers to the set of predicted probabilities. $P_{i,j}$ and $G_{i,j}$ represent the predicted probability and the one-hot encoded true value of class $j$ at voxel $i$, respectively. $I$ denotes the total number of voxels, and $J$ denotes the number of classes (Jiang et al., 2025).
7 Evaluation Metrics
We have adopted two primary evaluation metrics to assess segmentation performance: the Dice Similarity Coefficient (DSC), which quantifies the overlap between predicted and ground truth regions, and the 95th percentile Hausdorff Distance (HD95), which measures the spatial distance between boundary surfaces while mitigating the influence of outliers. Detailed definitions and computation procedures for these metrics are provided in the following subsection.
Hausdorff Distance (HD95)
The HD 95 is a boundary-based metric that evaluates segmentation quality by computing the 95th-percentile distance between the predicted volume’s boundary voxels and those of the ground-truth segmentation.
$HD_{95}(Y,P)\;=\;\max\bigl{(}d_{95}(Y,P),\,d_{95}(P,Y)\bigr{)}$ (9)
Here, $d_{95}(Y,P)$ is the maximum 95th percentile distance between the ground truth and predicted voxels, and $d_{95}(P,Y)$ is the maximum 95th percentile distance between the predicted and ground truth voxels.
Dice Similarity Coefficient (DSC).
The Dice Similarity Coefficient (DSC) measures the similarity between two sets, returning values from 0 to 1, with 1 representing perfect similarity. It is computed using the following formula:
$DSC(G,P)=\frac{2\,|G\cap P|}{|G|+|P|}=\frac{2\sum_{i=1}^{I}G_{i}P_{i}}{\sum_{i=1}^{I}G_{i}+\sum_{i=1}^{I}P_{i}}$ (10)
Where $G$ is the set of real results, $P$ refers to the set of predicted results, $G_{i}$ and $P_{i}$ represent the true and predicted values of the voxel i, respectively, and I is the number of voxels.
8 Results
In this section we summarise the results obtained on two different datasets. As previously said, we have compared the performance of our model with other convolution-based image segmentation models, including Swin-UNet (Cao et al., 2023), UNETR (Hatamizadeh et al., 2021), UNETR++ (Shaker et al., 2024), TransUNet (Chen et al., 2021), MISSFormer (Huang et al., 2023), nnFormer (Zhou et al., 2023), LeVit-UNet (Feng et al., 2024), PCCTrans (Xu et al., 2024) and others. A numerical and visual comparison among the results of the AMBER-AFNO model with these other modelsis performed.
8.1 ACDC Dataset
Tab. 2 and 3 show that AMBER–AFNO achieves the highest overall Dice score in the ACDC validation set (92. 85%), outperforming UNETR ++ (92. 83%) while using fewer than half of its parameters. The proposed model delivers the best myocardium segmentation (90.74%) and ranks second on both the right-ventricle (91.60%) and left-ventricle (96.21%) classes, demonstrating balanced accuracy across all cardiac structures. These results confirm that the AFNO encoder, combined with a lightweight SegFormer-style decoder, can match, or even surpass, state-of-the-art baselines in ACDC with substantially reduced computational demands.
Table 2: Dice Similarity Coefficient (%) on the ACDC validation set for the right ventricle (RV), myocardium (Myo) and left ventricle (LV); the column DSC reports the mean of the three structures. Bold values are the best in each column, while underlined values are the second best.
Methods RV Myo LV DSC
TransUNet(Chen et al., 2021) 88.86 84.54 95.73 89.71
Swin-UNet(Cao et al., 2023) 88.55 85.62 95.83 90.00
UNETR(Hatamizadeh et al., 2021) 85.29 86.52 94.02 86.61
MISSFormer(Huang et al., 2023) 86.36 85.75 91.59 87.90
nnFormer(Zhou et al., 2023) 90.94 89.58 95.65 92.06
UNETR++(Shaker et al., 2024) 91.89 90.61 96.00 92.83
LeVit-UNet(Feng et al., 2024) 89.55 87.64 93.76 90.32
PCCTrans(Xu et al., 2024) 90.55 90.57 96.22 92.45
AMBER–AFNO (ours) 91.60 90.74 96.21 92.85
Table 3: Comparison on ACDC. AMBER-AFNO achieves best segmentation results(DSC), while being efficient (Params in millions)
Methods Params FLOPs DSC
UNETR++(Shaker et al., 2024) 81.55 52.14 92.83
AMBER–AFNO (ours) 14.77 163.27 92.85
Table 4: Dice scores for eight abdominal organs and HD95 on the Synapse validation set. Bold values denote the best result in each column, while underlined values denote the second best.
Methods Spl RKid LKid Gal Liv Sto Aor Pan HD95 DSC
U-Net(Ronneberger et al., 2015) 86.67 68.60 77.77 69.72 93.43 75.58 89.07 53.98 76.85
TransUNet(Chen et al., 2021) 85.08 77.02 81.87 63.16 94.08 75.62 87.23 55.86 31.69 77.49
Swin-UNet(Cao et al., 2023) 90.66 79.61 83.28 66.53 94.29 76.60 85.47 56.58 21.55 79.13
UNETR(Hatamizadeh et al., 2021) 85.00 84.52 85.60 56.30 94.57 70.46 89.80 60.47 18.59 78.35
MISSFormer(Huang et al., 2023) 91.92 82.00 85.21 68.65 94.41 80.81 86.99 65.67 18.20 81.96
nnFormer(Zhou et al., 2023) 90.51 86.25 86.57 70.17 96.84 86.83 92.04 83.35 10.63 86.57
Swin-UNETR(Hatamizadeh et al., 2022) 95.37 86.26 86.99 66.54 95.72 77.01 91.12 68.80 10.55 83.48
UNETR++(Shaker et al., 2024) 95.77 87.18 87.54 71.25 96.42 86.01 92.52 81.10 7.53 87.22
LeVit-UNet(Feng et al., 2024) 88.86 80.25 84.61 62.23 93.11 72.76 87.33 59.07 16.84 78.53
PCCTrans(Xu et al., 2024) 88.84 82.64 85.49 68.79 93.45 71.88 86.59 66.31 17.10 80.50
AMBER–AFNO (ours) 87.82 86.26 87.36 61.33 96.02 80.50 91.42 79.36 16.96 83.76
Table 5: Comparison on Synapse. AMBER-AFNO achieves the third best segmentation results (DSC), while being more efficient (Params in millions)
Methods Params FLOPs DSC
TransUNet(Chen et al., 2021) 96.07 88.91 77.49
UNETR(Hatamizadeh et al., 2021) 92.49 75.76 78.35
Swin-UNet(Cao et al., 2023) 62.83 384.2 83.48
nnFormer(Zhou et al., 2023) 150.5 213.4 86.57
UNETR++(Shaker et al., 2024) 42.96 47.98 87.22
AMBER–AFNO (ours) 14.86 161.24 83.76
Figure 4: Visual Comparison of AMBER-AFNO and UNETR++ Model on ACDC Dataset
8.2 Synapse Dataset
[…]
8.3 Discussion: Performance vs Efficiency Trade-off
The experimental results reported in Tables 2, 3, 4 and 5 highlight the strong performance of the proposed AMBER-AFNO model when compared with both convolutional and transformer-based segmentation architectures. On the ACDC dataset, AMBER-AFNO achieves the highest overall Dice Similarity Coefficient (92.85%), slightly surpassing even the more complex UNETR++ model, which requires more than five times the number of parameters. This result is particularly significant as it confirms that frequency-domain mixing via AFNO can capture global contextual information without the computational burden of traditional self-attention.
On the Synapse dataset, which includes a greater anatomical variety and more heterogeneous imaging conditions, AMBER-AFNO remains highly competitive. While it is marginally outperformed by the heaviest models in terms of overall DSC, it still achieves a very solid performance (83.76%), only a few points behind state-of-the-art architectures like UNETR++ and nnFormer, but with a significantly lower parameter count and similar inference-time complexity. This reinforces the model’s capacity to generalize well across different domains and target structures without the need for extensive task-specific tuning.
The observed trade-off between performance and efficiency strongly supports the core motivation of this study. AMBER-AFNO demonstrates that it is possible to design architectures for 3D medical image segmentation that retain high representational power while dramatically reducing the number of trainable parameters. This is achieved through the use of Adaptive Fourier Neural Operators, which perform learnable filtering in the spectral domain and eliminate the quadratic scaling associated with attention mechanisms. Such a design makes AMBER-AFNO particularly suitable for deployment in resource-constrained environments, including clinical systems with limited memory and compute budgets.
Furthermore, the model achieves these results without relying on complex decoder branches, multi-scale fusion strategies, or deep attention hierarchies. The simplicity and modularity of the architecture make it easy to adapt to different datasets and segmentation tasks with minimal overhead. Taken together, these aspects make AMBER-AFNO not only a promising benchmark for lightweight segmentation, but also a practical choice for real-world applications where efficiency and robustness are equally important as accuracy.
Figure 5: Visual Comparison of AMBER-AFNO and UNETR++ Model on Synapse Dataset
[…]
9 Ablation Study
To isolate the performance attributable to Adaptive Fourier Neural Operators (AFNO), we train two encoder variants on the ACDC dataset using identical hyper-parameters—network depth, embedding width, number of heads, MLP expansion ratio, and deep-supervision strategy (cf. §3.1). The only difference lies in the attention mechanism.
  • AMBER (MHSA). A light configuration whose AFNO blocks are replaced by standard multi-head self-attention (MHSA); the embedded dimensions are:
    [23, 64, 128, 256][23,\,64,\,128,\,256][ 23 , 64 , 128 , 256 ].
  • AMBER-AFNO (light). Uses the AFNO block described in 3.1 with the same light configuration above.
[…]
Table 6: Efficiency–accuracy trade-off on the ACDC validation set. AFNO delivers state-of-the-art Dice scores with markedly fewer parameters. Best results are highlighted in bold, and second-best results are underlined.
Method Params FLOPs DSC (%)
UNETR++(Shaker et al., 2024) 81.55 52.14 92.83
AMBER-AFNO 14.77 163.27 92.85
AMBER-AFNO (light) 8.70 58.29 92.83
AMBER (MHSA) 19.01 132.07 92.03
10 Conclusions
In this study, we introduced AMBER-AFNO, a lightweight yet high-performing architecture for 3D medical image segmentation that replaces conventional attention mechanisms with Adaptive Fourier Neural Operators (AFNO). Originally developed for multiband remote sensing tasks, the AMBER framework was adapted to volumetric medical imaging, resulting in a model that achieves state-of-the-art performance with a fraction of the parameters required by leading transformer-based methods.
Through extensive experiments on two public benchmarks — ACDC and Synapse — we demonstrated that AMBER-AFNO achieves accuracy on par with or exceeding that of heavier architectures like UNETR++ and nnFormer, while reducing the parameter count by over 80%. This translates into faster training times, lower inference latency, and a reduced memory footprint—features especially valuable for clinical applications and edge deployment.
While AMBER-AFNO achieves top performance on the ACDC dataset, the results on the more heterogeneous Synapse benchmark are slightly behind the best-performing models. Current efforts are focused on investigating this performance gap, with the aim of further enhancing the model’s ability to generalize across diverse anatomical structures and imaging modalities.
Overall, this work underscores the promise of frequency-domain mixing in 3D medical imaging and highlights AMBER-AFNO as a compelling alternative for resource-efficient, high-accuracy segmentation.
Acknowledgments
This work has been funded by project code PIRO1_00011 “IBISCo”, PON 2014–2020, for all three entities (INFN, UNINA, and CNR).